Electric field strength due to dipole of moment p at a distant point r along the axis of the dipole is
1
Expert's answer
2019-07-15T09:25:10-0400


Electric field strength in point A

"E=\frac{q}{4\pi \epsilon_0}\left(\frac{1}{r_2^2}-\frac{1}{r_1^2}\right)"

We may apply the mean value theorem

"E=\frac{q}{4\pi \epsilon_0}\frac{d}{dr} \left(\frac{1}{r^2} \right) \bigg|_{r=\xi}(r_2-r_1)=\frac{1}{4\pi \epsilon_0}\frac{2p}{\xi^3}"

where

"\xi \in [r_2,r_1]"

Dipole is a point particle, hence

"\xi=r"

where r is a distance between dipole and point A.

So the electric field strength due to dipole of moment p at a distant point r along the axis of the dipole is

"E=\frac{1}{4\pi \epsilon_0}\frac{2p}{r^3}"

or in vector form

"\vec E=\frac{1}{4\pi \epsilon_0}\frac{2\vec p}{r^3}"